40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Sep 18 13:33:17 2015, ericq, Update, LSC, Fast ALS troubles - Noise at 36kHz IRlock_noises.pdf
    Reply  Fri Sep 18 16:08:41 2015, ericq, Update, LSC, Fast ALS troubles - Noise at 36kHz PRMIlock_REFLspectra.pdfIMCspectrum.pdf
       Reply  Fri Sep 18 19:15:35 2015, rana, Update, LSC, Fast ALS troubles - Noise at 36kHz 
    Reply  Tue Sep 29 16:52:49 2015, ericq, Update, LSC, Fast ALS troubles - unknown zero OLG_fastALS.pdf
       Reply  Tue Sep 29 18:03:11 2015, rana, Update, LSC, use LISO Screen_Shot_2015-09-29_at_5.55.34_PM.png
          Reply  Wed Sep 30 10:00:02 2015, ericq, Update, LSC, used LISO 
       Reply  Fri Oct 2 03:29:16 2015, ericq, Update, LSC, Fast ALS progress - AO path crossed over, but no high BW crossedover.pdf
          Reply  Fri Oct 9 13:31:21 2015, ericq, Update, LSC, ALS plant shape xarm_plantTFs.pdf
             Reply  Fri Oct 9 14:50:18 2015, Koji, Update, LSC, ALS plant shape 
                Reply  Fri Oct 9 16:23:25 2015, ericq, Update, LSC, ALS plant shape 
                   Reply  Tue Oct 13 16:28:21 2015, ericq, Update, LSC, Fast ALS pomona liso_lp.zip2LPfit.pdf
Message ID: 11620     Entry time: Fri Sep 18 13:33:17 2015     Reply to this: 11621   11648
Author: ericq 
Type: Update 
Category: LSC 
Subject: Fast ALS troubles - Noise at 36kHz 

To get around the problems between the pomona LPF and low CM board input impedance, I've placed the LPF at the CM board fast output. This won't work as a permanent solution, since we only want to lowpass the ALS signal, but it should be fine for a single arm test. 

However, I kept getting blown out of lock when turning up the AO gain, but well before I really expect any real action from the fast path. Looking at the OLTF, I was seeing some large spike at ~36kHz nearing 0dB loop gain with unstable phase. This prompted me to look at the ALS error signal out to higher bandwidth with the SR785; before I only ever looked at it through the digital system. 

So, with the X arm locked via POX11 I, and ITMY misaligned to use AS55 as an out of loop sensor, I measured the spectrum of the I ouput of the ALS X demod board (which was set to be near a zero crossing via the delay line), and the Q Mon of the AS55 demod board. 

Both ALS and AS55 show a sharp line at around 36.5kHz, so something is really happening in the IFO at this frequency. Koji might have seen an indication of this back in March.

What's going on here? And what would be different about PRFPMI that wouldn't have made this a problem for locking?

Attachment 1: IRlock_noises.pdf  35 kB  | Hide | Hide all
IRlock_noises.pdf
ELOG V3.1.3-